OpenStudy (anonymous):

A sequence is defined recursively using the formula f(n + 1) = –0.5 f(n) . If the first term of the sequence is 120, what is f(5)?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

helpp

OpenStudy (michele_laino):

please, note that, from data of your text, I can write this: \[f(2)=-\frac{ 1 }{ 2 }f(1)=-\frac{ 1 }{ 2 }*120=-60\]

OpenStudy (michele_laino):

\[f(3)=-\frac{ f(2) }{ 2 }=\frac{ -60 }{ 2 }=-30\]

OpenStudy (anonymous):

the way i put it is the right way there is no difference

OpenStudy (michele_laino):

oops..f(39=-(-60/2)=30

OpenStudy (michele_laino):

f(3)=30

OpenStudy (michele_laino):

\[f(4)=-\frac{ f(3) }{ 2 }=-\frac{ 30 }{ 2 }=-15\]

OpenStudy (michele_laino):

finally: \[f(5)=-\frac{ f(4) }{ 2 }=-\frac{ -15 }{ 2 }=\frac{ 15 }{ 2 }\]

OpenStudy (anonymous):

another one plssss

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

we have: \[f(290f(1)-8=100-8=92\]

OpenStudy (michele_laino):

oops..f(2)=f(1)-8=100-8=92

OpenStudy (michele_laino):

f(3)=f(2)-8=92-8=84

OpenStudy (michele_laino):

f(4)=f(3)-8=84-8=76

OpenStudy (michele_laino):

f(5)=f(4)-8=76-8=68

OpenStudy (michele_laino):

finally: f(6)=f(5)-8=68-8=... please complete the computation

OpenStudy (anonymous):

60

OpenStudy (michele_laino):

that's right!